If I have 10,000 dollars to invest at 8% compounded quarterly, how long will it take to earn $5,000 in interest. Please can you provide a formula for this problem.
Thank you

Answer by scott8148(6628)

the formula is ___ A = P (1 + r)^n
___ A = current Amount
___ P = starting Principal
___ r = interest rate per compounding period
___ n = number of compounding periods

in this case ___ 15000 = 10000 (1 + .02)^(t/4) (where t is in years)

dividing by 10000 ___ 1.5 = (1.02)^(t/4)

taking logarithm ___ log(1.5) = (t/4) log(1.02)

dividing by log(1.02) ___ [log(1.5)] / [log(1.02)] = t / 4

multiplying by 4 ___ 4 {[log(1.5)] / [log(1.02)]} = t